a group of 6 men and 4 women.... three people are selected to attend a conference.
in how man ways can group of 3 women be selected from the 4
if all groups are equally likely to be selected whats the probability that the group will conisist of all women

Answer by stanbon(75887)

a group of 6 men and 4 women.... three people are selected to attend a conference.
In how many ways can a group of 3 women be selected from the 4
4C3 = 4C1 = 4
-----------------------
If all groups are equally likely to be selected?
Whats the probability that the group will consist of all women
---
P(all women) = 4/10C3 = 4/120
================================
Cheers,
Stan H.